# Temporal Analysis and Repair of Flaky Dockerfiles

### v1

@@ VII Related Work
Studies on Dockerfiles. In Sections I and II, we presented the most recent works on Dockerfile analysis [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. FlakiDock differentiates itself from existing Dockerfile smell detection tools [2, 9, 10, 7] and repair tools such as [4] by not only addressing static issues within Dockerfiles but also targeting the dynamic errors caused by the flakiness of Dockerfiles. Unlike existing tools that primarily rely on pre-defined static analysis or require manual interventions while analyzing dynamic information [6], FlakiDock uses LLMs and retrieval-augmented techniques, automatically analyzing both static and dynamic information. Furthermore, FlakiDock employs a feedback loop to incorporate false repairs to alleviate LLM mistakes.
@@
VIII Conclusion
This paper presents the first comprehensive study on Dockerfile flakiness, revealing that 9.81% of Dockerfiles exhibit flaky behavior, which impacts the reliability of CI/CD pipelines. We introduce the first taxonomy of Dockerfile flakiness and propose FlakiDock, a novel tool that leverages large language models, retrieval-augmented generation, dynamic analysis, and an iterative feedback loop for automatic Dockerfile flakiness repair. Our evaluation shows that FlakiDock achieves a 73.55% repair accuracy, outperforming existing tools like Parfum by 12,581% and GPT-4 based prompting by 94.63%. These results highlight the effectiveness of FlakiDock in addressing Dockerfile flakiness. As part of our future work, we plan to extend FlakiDock to handle even more intricate build scenarios.
IX Data Availability
We have made our dataset, model, comparison framework, and FlakiDock’s implementation available [15] for the reproducibility of results. We further provide instructions for replicating our experimental setup.

### v2

@@ VIII Related Work
Dockerfile Analysis. In Sections I and II, we discussed recent work on Dockerfile analysis [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. FlakiDock differentiates itself from existing Dockerfile smell detection tools [3, 10, 11, 8] and repair tools such as [5] by not only addressing static issues within Dockerfiles but also targeting the dynamic errors caused by the flakiness of Dockerfiles. Shipwright [7] introduces a human-in-the-loop approach for repairing broken Dockerfiles. Shipwright begins by clustering broken Dockerfiles to identify common failure patterns. This is achieved by embedding build log text into vector representations using a modified BERT language model, followed by clustering with the HDBSCAN [13] algorithm. Representative Dockerfiles from each cluster are then presented to a human supervisor, who searches for solutions to the identified failures. Discovered solutions are evaluated for applicability across the cluster, and successful fixes are formalized into a database of repair patterns. Each pattern comprises a regular expression to identify issues and a corresponding repair function detailing the fix. In contrast, FlakiDock is fully automated, leveraging LLMs, retrieval-augmented techniques, and a feedback loop to refine repairs without requiring manual oversight.
@@
IX Conclusion
We present the first comprehensive study on Dockerfile flakiness, identifying that 9.81% of Dockerfiles exhibit flaky behavior, significantly impacting the reliability of CI/CD pipelines. To address this issue, we introduce a novel taxonomy of Dockerfile flakiness and propose FlakiDock that leverages language models, retrieval-augmented generation, dynamic analysis, and an iterative feedback loop for automatic repair. Our evaluation shows that FlakiDock achieves a 73.55% repair accuracy, outperforming Parfum, Shipwright, and GPT-4-based prompting by 12,581%, 1,232%, and 95%, respectively. These results highlight the effectiveness of FlakiDock in addressing Dockerfile flakiness.
X Data Availability
We have made our dataset, model, comparison framework, and FlakiDock’s implementation available [20] for the reproducibility of results.